# Tailored Conversations beyond LLMs: A RL-Based Dialogue Manager

### v1

@@ 1 Introduction
In recent years, the demand for mental health services has surged, outpacing the availability of resources and creating a substantial gap in access to care Cameron et al. (2017). As a result, many patients face extended waiting periods before receiving therapy Cameron et al. (2017); Denecke et al. (2020). To address this challenge, virtual agents capable of simulating Motivational Interviewing (MI) have been proposed as a potential interim support system for individuals awaiting treatment. These agents can provide immediate assistance, particularly beneficial in therapeutic approaches requiring multiple sessions Fiske et al. (2019). However, their role is not to replace human therapists but to serve as a supplementary tool that enhances existing therapeutic interventions. Motivational Interviewing (MI) poses a particularly complex challenge for dialogue systems, traditionally addressed through intricate rule-based frameworks Prochaska et al. (2021); Olafsson et al. (2020). However, recent advances in natural language processing (NLP) have paved the way for leveraging large language models (LLMs) such as GPT-like architectures Baktash and Dawodi (2023) in such applications Steenstra et al. (2024), significantly expanding the scope of dialogue systems across various domains. Although these models exhibit remarkable language generation capabilities, they also present significant limitations, many of which can be addressed through insights from "traditional" dialogue research. In particular, LLMs often lack the controllability and structured decision-making of conventional rule-based systems, which are more predictable and interpretable Shidara et al. (2020). Rule-based domain-specific dialogue systems Hadi et al. (2024) offer notable advantages, including improved controllability, explainability, and the ability to integrate expert knowledge. However, they are typically less adaptable and more resource-intensive to develop. In contrast, LLMs demonstrate strong adaptability across domains but pose challenges in achieving control, transparency, and efficiency. Additionally, incorporating expert knowledge into LLMs often requires extensive domain-specific data Hadi et al. (2024). Notably, reinforcement learning (RL)-based dialogue managers Pecune and Marsella (2020) have shown promise in enhancing control and coherence in dialogue systems. Hence, a promising approach involves hybrid models that combine the strengths of both paradigms—leveraging the adaptability and generative capabilities of LLMs while integrating a domain-specific dialogue manager to regulate interactions Abu-Rasheed et al. (2024); Galland et al. (2024).
@@ 2.1 Patients’ Profiles in MI
Patients participating in motivational interviewing (MI) exhibit varying levels of readiness to change their behaviors. As proposed by anonymous (2024a), these patients can be classified into three categories: Open-to-Change, Resistant-to-Change, and Receptive. Open-to-Change individuals demonstrate a strong willingness to modify unhealthy behaviors. Resistant-to-Change patients are generally reluctant to alter their current behaviors, showing a preference for maintaining the status quo. Receptive patients, while initially exhibiting low motivation to change, gradually develop a higher motivation to adopt healthier behaviors as the conversation progresses. These classifications capture variations in patients’ responses and therapists’ strategies, as discussed in anonymous (2024a). The ability to adapt the flow of dialogue to these three patient profiles can significantly enhance the efficiency of the therapist’s dialogue model. A dialogue system for MI should be able to take into account the particular challenges that such dialogues pose. Such a dialogue system should be able to navigate across phases while being able to adapt to different profiles of users.
@@ 3 Related Work
Motivational Interviewing (MI) presents significant challenges for dialogue systems, as it necessitates both a structured progression through its four distinct phases —engagement, focusing, evoking, and planning Miller and Rollnick (2012) — and adaptability to diverse patients’ profiles anonymous (2024a). While existing systems such as Woebot Prochaska et al. (2021) have demonstrated the feasibility of MI-based chatbots by incorporating therapeutic frameworks like cognitive behavioral therapy (CBT) and mindfulness, they predominantly rely on static or rule-based architectures. Steenstral et al. Steenstra et al. (2024) identified the limitations of rule-based approaches in maintaining adherence to therapeutic protocols and proposed leveraging LLMs for this application, demonstrating promising results. The advent of large language models (LLMs) has transformed dialogue generation, offering new possibilities for MI-based interactions Steenstra et al. (2024). Models such as GPT-like systems Baktash and Dawodi (2023) exhibit strong generative capabilities and adaptability across diverse applications, which can be leveraged to enhance the social dimensions of interaction. This aspect is particularly relevant, as Kanaoka et al. Kanaoka and Mutlu (2015) emphasized the critical role of social engagement and rapport in facilitating behavioral change. Recent research has underscored the importance of cognitive modeling and adaptability in dialogue systems to more accurately account for the mental states of both the agent and the user. For instance, He et al. He et al. (2024) introduced dual reasoning mechanisms that enable LLMs to incorporate contextual nuances, while Zhang et al. Zhang et al. (2020) explored interactive agent representations to improve dialogue coherence. Despite these advancements, LLMs remain constrained in terms of controllability and domain specificity Shidara et al. (2020). Recent efforts have sought to enhance control and applicability, particularly within task-oriented dialogue systems. Yao et al. Yao et al. (2023) employed reinforcement learning (RL) to optimize LLM prompting strategies; however, black-box LLMs continue to exhibit limitations in controllability and often generate repetitive responses. To address these shortcomings, Xu et al. Xu et al. (2023) demonstrated the advantages of integrating fine-tuned, smaller language models with larger LLMs, while Yu et al. Yu et al. (2023) employed Monte Carlo tree search for optimal action selection, improving both coherence and practical utility. Although these studies mark significant progress in hybrid model design, they often fall short of accommodating a broad spectrum of user profiles. Integrating the structured controllability of classical dialogue models with the generative flexibility of LLMs within an adaptive hybrid framework could facilitate more dynamic, personalized, and effective MI interactions. Reinforcement learning has been instrumental in optimizing dialogue policies and refining system behavior. Traditional RL-based approaches have primarily focused on enhancing user engagement and task success rates. Walker et al. Walker (2000) and Li et al. Li et al. (2016) demonstrated the efficacy of RL in training conversational agents for goal-oriented tasks, while Weber et al. Weber et al. (2018) illustrated its effectiveness in selecting contextually appropriate actions, such as humor or sound effects, to enrich user experiences. More advanced frameworks have incorporated both social and task-oriented rewards Pecune and Marsella (2020) or jointly trained user and dialogue policies Takanobu et al. (2020). Although these methods enable adaptation to different user profiles, they remain largely confined to task-oriented systems that rely on predefined natural language templates. In this work, we propose a novel framework that integrates LLMs with an RL-based dialogue manager to structure MI dialogues across different phases while dynamically adapting to diverse patient profiles. By synergizing structured control with generative flexibility, our approach enhances adaptability and efficiency, enabling the system to learn from limited data, transit fluidly between MI phases, and personalize responses to heterogeneous patient needs. The remainder of this paper is organized as follows: Section 4 details our proposed methodology, while Section 5 describes our evaluation environment. Section 6 presents the experimental results, and Section 7 provides an analysis and interpretation of our findings.
@@ 4.1 Problem description
The objective of the proposed model is to predict the optimal action $a_{t}$ at each time step $t$, given the dialogue context $c$. Each action corresponds to a dialogue act representing the virtual therapist’s strategic behavior. The agent sentence is then generated by a conditioned large language model (LLM) that produces an utterance coherent with the context and realizes the selected dialogue act, as validated in anonymous (2024b).
@@ 5.2.1 Simulate patients in MI
Simulating patients in MI has been explored in prior research. For instance, anonymous (2024c) proposed a prompt to simulate a user with a large language model (LLM) anonymous (2024b). This approach has been validated to produce contextually relevant, natural dialogue acts and utterances anonymous (2024b, c), although the differences between user profiles have not been tested and might be a limitation of this simulator. We use this user simulator to train and evaluate our dialogue manager. In the following of this article, the term user refers to this user simulator.
@@ 5.2.2 Action Space
The agent operates in a discrete action space consisting of 13 possible dialogue acts, which are categorized into task-oriented dialogue acts and socially oriented dialogue acts. Task-oriented dialogue acts include Asking for Consent or Validation, Providing Medical Education and Guidance, Planning with the Patient, Giving a Solution, Asking about Current Emotions, Inviting a Shift in Outlook, Asking for Information, and Reflection. Socially-oriented dialogue acts include Empathic reactions, Acknowledging Progress and Encouraging, Backchanneling, Greeting or Closing, and Normalizing Experiences while Providing Reassurance. This taxonomy was introduced in anonymous (2024a).
@@ 7 Interpretation
In this section, we examine the impact of meta-learning on training the master policy. Figure 4 illustrates the distribution of master actions activation over time for different user profiles for the full model. For the model trained without meta-learning, we observe a collapse of the master policy to a single dominant action across interactions. This highlights the difficulty of learning a generalized policy that performs well across diverse user profiles without explicit mechanisms for adaptation. In contrast, the use of meta-learning allows the master policy to maintain variability and adaptability in its actions. In Figure 4, we associate each master action with MI phases based on the predominant dialogue acts typically observed in those phases and master actions, providing interpretability into the policy’s learned structure. The models effectively differentiate between distinct phases, initially engaging in Engaging/Focusing phases (Master action 5) before transitioning into Focusing/Planning phases (Master action 3) and Evoking phase (Master action 4).The engaging phase is quicker for Open to change users, as they do human therapist, aligning with prior findings anonymous (2024a). Indeed, open-to-change patients require less rapport-building with the therapist, as they choose to initiate the conversation. This qualitative analysis highlights the benefits of incorporating meta-learning, as it enhances the model’s ability to structure MI phases effectively and provide a more personalized dialogue experience.

### v2

@@ 1 Introduction
In recent years, the demand for mental health services has surged, outpacing the availability of resources and creating a substantial gap in access to care Cameron et al. (2017). As a result, many patients face extended waiting periods before receiving therapy Cameron et al. (2017); Denecke et al. (2020). To address this challenge, virtual agents capable of simulating Motivational Interviewing (MI) have been proposed as a potential interim support system for individuals awaiting treatment. These agents can provide immediate assistance, particularly beneficial in therapeutic approaches requiring multiple sessions Fiske et al. (2019). However, their role is not to replace human therapists but to serve as a supplementary tool that enhances existing therapeutic interventions. Motivational Interviewing (MI) poses a particularly complex challenge for dialogue systems, traditionally addressed through intricate rule-based frameworks Prochaska et al. (2021); Olafsson et al. (2020). However, recent advances in natural language processing (NLP) have paved the way for leveraging large language models (LLMs) such as GPT-like architectures Baktash and Dawodi (2023) in such applications Steenstra et al. (2024), significantly expanding the scope of dialogue systems across various domains. Although these models exhibit remarkable language generation capabilities, they also present significant limitations, many of which can be addressed through insights from "traditional" dialogue research. In particular, LLMs often lack the controllability and structured decision-making of conventional rule-based systems, which are more predictable and interpretable Shidara et al. (2020). Rule-based domain-specific dialogue systems Hadi et al. (2024) offer notable advantages, including improved controllability, explainability, and the ability to integrate expert knowledge. However, they are typically less adaptable and more resource-intensive to develop. In contrast, LLMs demonstrate strong adaptability across domains but pose challenges in achieving control, transparency, and efficiency. Additionally, incorporating expert knowledge into LLMs often requires extensive domain-specific data Hadi et al. (2024). Notably, reinforcement learning (RL)-based dialogue managers Pecune and Marsella (2020) have shown promise in enhancing control and coherence in dialogue systems. Hence, a promising approach involves hybrid models that combine the strengths of both paradigms—leveraging the adaptability and generative capabilities of LLMs while integrating a domain-specific dialogue manager to regulate interactions Abu-Rasheed et al. (2024); Galland et al. (2024b).
@@ 2.1 Patients’ Profiles in MI
Patients participating in motivational interviewing (MI) exhibit varying levels of readiness to change their behaviors. As proposed by anonymous (2024), these patients can be classified into three categories: Open-to-Change, Resistant-to-Change, and Receptive. Open-to-Change individuals demonstrate a strong willingness to modify unhealthy behaviors. Resistant-to-Change patients are generally reluctant to alter their current behaviors, showing a preference for maintaining the status quo. Receptive patients, while initially exhibiting low motivation to change, gradually develop a higher motivation to adopt healthier behaviors as the conversation progresses. These classifications capture variations in patients’ responses and therapists’ strategies, as discussed in anonymous (2024). The ability to adapt the flow of dialogue to these three patient profiles can significantly enhance the efficiency of the therapist’s dialogue model. A dialogue system for MI should be able to take into account the particular challenges that such dialogues pose. Such a dialogue system should be able to navigate across phases while being able to adapt to different profiles of users.
@@ 3 Related Work
Motivational Interviewing (MI) presents significant challenges for dialogue systems, as it necessitates both a structured progression through its four distinct phases —engagement, focusing, evoking, and planning Miller and Rollnick (2012) — and adaptability to diverse patients’ profiles anonymous (2024). While existing systems such as Woebot Prochaska et al. (2021) have demonstrated the feasibility of MI-based chatbots by incorporating therapeutic frameworks like cognitive behavioral therapy (CBT) and mindfulness, they predominantly rely on static or rule-based architectures. Steenstral et al. Steenstra et al. (2024) identified the limitations of rule-based approaches in maintaining adherence to therapeutic protocols and proposed leveraging LLMs for this application, demonstrating promising results. The advent of large language models (LLMs) has transformed dialogue generation, offering new possibilities for MI-based interactions Steenstra et al. (2024). Models such as GPT-like systems Baktash and Dawodi (2023) exhibit strong generative capabilities and adaptability across diverse applications, which can be leveraged to enhance the social dimensions of interaction. This aspect is particularly relevant, as Kanaoka et al. Kanaoka and Mutlu (2015) emphasized the critical role of social engagement and rapport in facilitating behavioral change. Recent research has underscored the importance of cognitive modeling and adaptability in dialogue systems to more accurately account for the mental states of both the agent and the user. For instance, He et al. He et al. (2024) introduced dual reasoning mechanisms that enable LLMs to incorporate contextual nuances, while Zhang et al. Zhang et al. (2020) explored interactive agent representations to improve dialogue coherence. Despite these advancements, LLMs remain constrained in terms of controllability and domain specificity Shidara et al. (2020). Recent efforts have sought to enhance control and applicability, particularly within task-oriented dialogue systems. Yao et al. Yao et al. (2023) employed reinforcement learning (RL) to optimize LLM prompting strategies; however, black-box LLMs continue to exhibit limitations in controllability and often generate repetitive responses. To address these shortcomings, Xu et al. Xu et al. (2023) demonstrated the advantages of integrating fine-tuned, smaller language models with larger LLMs, while Yu et al. Yu et al. (2023) employed Monte Carlo tree search for optimal action selection, improving both coherence and practical utility. Although these studies mark significant progress in hybrid model design, they often fall short of accommodating a broad spectrum of user profiles. Integrating the structured controllability of classical dialogue models with the generative flexibility of LLMs within an adaptive hybrid framework could facilitate more dynamic, personalized, and effective MI interactions. Reinforcement learning has been instrumental in optimizing dialogue policies and refining system behavior. Traditional RL-based approaches have primarily focused on enhancing user engagement and task success rates. Walker et al. Walker (2000) and Li et al. Li et al. (2016) demonstrated the efficacy of RL in training conversational agents for goal-oriented tasks, while Weber et al. Weber et al. (2018) illustrated its effectiveness in selecting contextually appropriate actions, such as humor or sound effects, to enrich user experiences. More advanced frameworks have incorporated both social and task-oriented rewards Pecune and Marsella (2020) or jointly trained user and dialogue policies Takanobu et al. (2020). Although these methods enable adaptation to different user profiles, they remain largely confined to task-oriented systems that rely on predefined natural language templates. In this work, we propose a novel framework that integrates LLMs with an RL-based dialogue manager to structure MI dialogues across different phases while dynamically adapting to diverse patient profiles. By synergizing structured control with generative flexibility, our approach enhances adaptability and efficiency, enabling the system to learn from limited data, transit fluidly between MI phases, and personalize responses to heterogeneous patient needs. The remainder of this paper is organized as follows: Section 4 details our proposed methodology, while Section 5 describes our evaluation environment. Section 6 presents the experimental results, and Section 7 provides an analysis and interpretation of our findings.
@@ 4.1 Problem description
The objective of the proposed model is to predict the optimal action $a_{t}$ at each time step $t$, given the dialogue context $c$. Each action corresponds to a dialogue act representing the virtual therapist’s strategic behavior. The agent sentence is then generated by a conditioned large language model (LLM) that produces an utterance coherent with the context and realizes the selected dialogue act, as validated in Galland et al. (2024b).
@@ 5.2.1 Simulate patients in MI
Simulating patients in MI has been explored in prior research. For instance, Galland et al. (2024a) proposed a prompt to simulate a user with a large language model (LLM) Galland et al. (2024b). This approach has been validated to produce contextually relevant, natural dialogue acts and utterances Galland et al. (2024b, a), although the differences between user profiles have not been tested and might be a limitation of this simulator. We use this user simulator to train and evaluate our dialogue manager. In the following of this article, the term user refers to this user simulator.
@@ 5.2.2 Action Space
The agent operates in a discrete action space consisting of 13 possible dialogue acts, which are categorized into task-oriented dialogue acts and socially oriented dialogue acts. Task-oriented dialogue acts include Asking for Consent or Validation, Providing Medical Education and Guidance, Planning with the Patient, Giving a Solution, Asking about Current Emotions, Inviting a Shift in Outlook, Asking for Information, and Reflection. Socially-oriented dialogue acts include Empathic reactions, Acknowledging Progress and Encouraging, Backchanneling, Greeting or Closing, and Normalizing Experiences while Providing Reassurance. This taxonomy was introduced in anonymous (2024).
@@ 7 Interpretation
In this section, we examine the impact of meta-learning on training the master policy. Figure 4 illustrates the distribution of master actions activation over time for different user profiles for the full model. For the model trained without meta-learning, we observe a collapse of the master policy to a single dominant action across interactions. This highlights the difficulty of learning a generalized policy that performs well across diverse user profiles without explicit mechanisms for adaptation. In contrast, the use of meta-learning allows the master policy to maintain variability and adaptability in its actions. In Figure 4, we associate each master action with MI phases based on the predominant dialogue acts typically observed in those phases and master actions, providing interpretability into the policy’s learned structure. The models effectively differentiate between distinct phases, initially engaging in Engaging/Focusing phases (Master action 5) before transitioning into Focusing/Planning phases (Master action 3) and Evoking phase (Master action 4).The engaging phase is quicker for Open to change users, as they do human therapist, aligning with prior findings anonymous (2024). Indeed, open-to-change patients require less rapport-building with the therapist, as they choose to initiate the conversation. This qualitative analysis highlights the benefits of incorporating meta-learning, as it enhances the model’s ability to structure MI phases effectively and provide a more personalized dialogue experience.